Three dice are rolled together. What is the probability as getting at least one '4'


1
Expert's answer
2021-10-19T10:19:52-0400

total number of ways:

"n=6^3=216"


number of ways to get one '4':

"n_1=3\cdot6^2=108"


number of ways to get two '4':

"n_2=3\cdot6=18"


number of ways to get three '4':

"n_3=1"


the probability of getting at least one '4'


"p=\frac{n_1+n_2+n_2}{n}=\frac{108+18+1}{216}=\frac{127}{216}"